OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6. y = 1/3x + 14 thirds y = 3x + 10 y = -3x - 2 y = -1/3x + 10 thirds

OpenStudy (anonymous):

parallel lines have the same slope. so, we are looking for an equation like this: y=-3x+b we can use the point (-2,4) and plug it in so the equation would be: (4)=-3(-2)+b (4)=6+b -2=b so, the new equation is y=-3x-2

OpenStudy (anonymous):

THANK YOU!!!!!!

OpenStudy (anonymous):

What if they were perpendicular?

OpenStudy (anonymous):

perpendicular lines have reciprocal opposite slopes, so the slope of a line perp. to a line with a slope of 2 would be -(1/2)

OpenStudy (anonymous):

a line perpendicular to \[y=-\frac{ 1 }{ 3 }x\] would have a slope of 3

OpenStudy (anonymous):

Oh ya I forgot about that!! So if a perpendicular line is y=-5/4 x it would be 4/5?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thank You!!!!! ={]